OpenStudy (anonymous):

Identify the solutions of 3x2 + 14x + 8 = 0.

OpenStudy (anonymous):

\[ (4+x) (2+3 x) \]

OpenStudy (anonymous):

x= -4 and x = -2/3

OpenStudy (anonymous):

now my assistant myininaya will probably show you the intermediate steps :P

myininaya (myininaya):

Well I like to find factors of a*c that have product a*c that add up to be b a=3,b=14,c=8 a*c=3(8)=24=12(2) b=14=12+2 bx=12x+2x so we can replace 14x with 12x+2x and then we will factor by grouping 3x^2+12x+2x+8=0 3x(x+4)+2(x+4)=0 (x+4)(3x+2)=0